-2/3 × 3/5 + 5/2 – 3/5 × 1/6
 -2/3 × 3/5 + 5/2 – 3/5 × 1/6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

-2/3 × 3/5 + 5/2–3/5 × 1/6
= -2/3 × 3/5–3/5 × 1/6+ 5/2 (by commutativity)​
= 3/5 (-2/3–1/6)+ 5/2
= 3/5 ((- 4–1)/6)+ 5/2
= 3/5 ((–5)/6)+ 5/2 (by distributivity)
= -15/30+5/2
= -1/2+5/2
= 4/2
= 2

​Thanks